OpenStudy (kathydimieri12):

Dona used a probability simulator to pull 3 colored marbles from a bag and flip a coin 50 times. The results are shown in the tables below: Color of Marble Number of Times Rolled Blue 22 Green 16 Yellow 12 Heads Tails 22 28 Using Dona's simulation, what is the probability of pulling a blue marble and the coin landing tails up? 512 over 2500 616 over 2500 50 over 50 68 over 50

OpenStudy (kathydimieri12):

@Mehek14 @kimberlyevens @inkyvoyd @OregonDuck

oregonduck (oregonduck):

Looks like it is 440/2500 = 0.176

OpenStudy (mathstudent55):

There were 22 blue marbles drawn out of 50 drawings. P(blue marble) = 22/50 There were 28 tails tossed out of 50 tosses P(tails) = 28/50 To find the overall probability of two independent events, multiply the 2 probabilities together. P(blue marble followed by tails) = P(blue marble) * P(tails)